OpenStudy (anonymous):

A hydrogen-filled balloon was ignited and 1.60g of hydrogen reacted with 12.8g of oxygen. How many grams of water vapor were formed? (Assume that water vapor is the only product.)

OpenStudy (noelgreco):

Conservation of mass.

OpenStudy (anonymous):

2H2+O2>>2H2O First we find the numbers of mole of each reactants to determine whether there is any limiting reagent. (The amount of product is affected by the limiting reagent in the reaction.) so using the equation numbers of mole= mass of reactant/ molar mass of reactant: n=m/M n(H2)=1.6/1.008 n(O2)=12.8/16 =1.587mol =0.8 mol from the equation, we know that 2 moles of H2 is needed to react with one mole of O2 so for 0.8 mol of O2, 1.6 mol of H2 is needed. So H2 is the limiting reagent. Again from the equation: 2 mol of H2 forms 2 mol of H2O so n(H2) = n(H2O) =1.587mol m(H2O)=nM =1.587 x (2x1.008+16.00) =28.6g